### v1

# KeepKV: Eliminating Output Perturbation in KV Cache Compression for Efficient LLMs Inference

Abstract
Efficient inference of large language models (LLMs) is hindered by an ever-growing key-value (KV) cache, making KV cache compression a critical research direction. Traditional methods selectively evict less important KV cache entries based on attention scores or position heuristics, which leads to information loss and hallucinations. Recently, merging-based strategies have been explored to retain more information by merging KV pairs that would be discarded; however, these existing approaches inevitably introduce inconsistencies in attention distributions before and after merging, causing output perturbation and degraded generation quality. To overcome this challenge, we propose KeepKV, a novel adaptive KV cache merging method designed to eliminate output perturbation while preserving performance under strict memory constraints. KeepKV introduces the Electoral Votes mechanism that records merging history and adaptively adjusts attention scores. Moreover, it further leverages a novel Zero Inference-Perturbation Merging methods, keeping attention consistency and compensating for attention loss resulting from cache merging. KeepKV successfully retains essential context information within a significantly compressed cache. Extensive experiments on various benchmarks and LLM architectures demonstrate that KeepKV substantially reduces memory usage, enhances inference throughput by more than 2×2$\times$2 × and keeps superior generation quality even with $10\%$ KV cache budgets.
1 Introduction
Transformer-based large language models (LLMs) have demonstrated remarkable capabilities across various applications [1, 2, 3, 4, 5, 6, 7]. To accelerate inference, LLMs commonly employ a key-value (KV) cache mechanism, which stores the KV embeddings of previously processed tokens to avoid redundant computations [8, 9, 10]. However, as LLMs continue to support increasingly longer context lengths, the size of the KV cache grows rapidly, becoming a major bottleneck for inference [11]. For example, in the case of LLaMA-3-70B, a batch size of 128 with an 8K context length requires up to 320GB of KV cache memory [2]. Consequently, compressing the KV cache while preserving generation quality has become a crucial challenge.
KV cache eviction is an effective approach to alleviating memory pressure. Existing studies primarily focus on selectively retaining important cache entries based on attention scores and token positions [12, 13, 14, 15, 16], as well as dynamically allocating varying cache sizes across layers and heads [17, 18, 19, 20]. However, eviction is irreversible—once a KV entry is evicted, its information is permanently lost, leading to issues such as hallucinations and context inconsistency [21]. To mitigate these drawbacks, recent studies have explored merging techniques [21, 22, 23, 24], which attempt to integrate cache entries that would otherwise be discarded into the remaining cache. CaM [21] first proposed merging the evicted value states into others. Recent methods such as D2O [23] and KVMerger [24] have proposed synchronized weighted merging strategies for both keys and values, achieving certain improvements. However, KV cache merging remains in an exploratory stage, particularly regarding merge candidate selection and merging weight computation. Consequently, existing strategies vary widely and lack solid theoretical foundations. For instance, D2O computes merging weights using cosine similarity, whereas KVMerger employs Gaussian kernel weights. We observe that existing methods inevitably introduce output perturbation. Specifically, the merged KV pair’s attention score is lower than the sum of the original scores prior to merging—a phenomenon we term "Attention Sag", as shown in Figure 1. These issues collectively underline the necessity for an efficient and theoretically grounded KV cache merging strategy.
Figure 1: Illustration of KeepKV vs. Existing Methods. The three middle blocks represent KV subject to eviction/merging. (a) Eviction methods permanently discard them. (b) Merging methods integrates them into retained KV, but the result is not equivalent to the full KV, causing "Attention Sag." (c) Full KV serves as the ideal baseline. (d) KeepKV uses Electoral Votes as merging records and applies ZIP-Merging to minimize output disturbance, ensuring consistency and improving performance.
To address this issue, we propose KeepKV, a novel merging method that maintains merging information and adaptively adjusts weights to eliminate output perturbation. The motivation behind KeepKV is to preserve the attention distribution and output consistency before and after merging at the current step, thus preventing any disturbances to inference. To achieve this, we first analyze the limitations of existing weighted merging methods in the context of attention computation, and propose the Electoral Votes mechanism, which records the number of times each KV pair has been merged, enabling the compressed cache to function equivalently to the original. Subsequently, we introduce Zero Inference-Perturbation Merging (ZIP-Merging), which automatically adjusts weights to compensate for any losses caused by merging, maintaining attention consistency. These designs theoretically guarantee zero attention-output perturbation at the current iteration despite compression.
Figure 2: Relative Performance on a summarization task at $5\%$ compression rate. KeepKV is closest to full KV ($100\%$).
To extend KeepKV to multi-step generation and practical tasks, we replace the current iteration’s attention scores in KeepKV computations with estimations based on recent historical scores. This is motivated by our empirical observation of strong locality in attention scores, also confirmed in prior studies [21, 25]. Crucially, we provide theoretical analyses guaranteeing bounded output perturbation across multiple steps, achieving a controllable upper bound. Moreover, we offer a theoretical interpretation for prevalent similarity-based candidate selection methods, incorporating it into our design. Through theoretical derivation and extensive experiments, we demonstrate that KeepKV effectively preserves attention stability and output consistency, outperforming state-of-the-art KV cache eviction and merging methods. Across various benchmark tasks, KeepKV achieves superior performance and is able to maintain strong results even with $5\%$ KV cache budgets, as shown in Figure 2. The contributions of this paper are summarized as follows:
  • We propose KeepKV, a novel adaptive KV cache merging approach designed to eliminate output perturbation caused by compression. KeepKV introduces the Electoral Votes mechanism and Zero Inference-Perturbation Merging to keep attention consistency.
  • Extensive experiments across various tasks and models show that KeepKV maintains better performance under limited cache, outperforming existing KV cache eviction and merging methods.
  • We are the first to theoretically analyze KV merging from the perspective of eliminating output perturbation. We provide guarantees on the perturbation bound of KeepKV and reveal the theoretical basis for merge candidate selection and weight design. Hopefully, our study can inspire future research on KV cache compression.
2 Related Work
KV cache has become a major bottleneck for efficient LLMs inference. Post-training optimization serves as a key solution due to its real-time and extensible capabilities.[26]. Existing methods fall into three categories: quantization, eviction, and merging.
KV Cache Quantization. Quantization methods convert tensor values to lower precision to reduce bit-width. KVQuant [27] applies Per-Channel Quantization for keys and Per-Token Quantization for values. MiKV [28] introduces mixed-precision KV caching, where less critical KV are stored at lower precision. Additionally, GEAR [29] leverages low-rank matrix approximation for quantization residuals to minimize quantization loss. Our KeepKVreduces the number of cached KV pairs through merging, which is orthogonal to quantization methods and can be combined for better efficiency.
KV Cache Eviction. Eviction methods only retain more important KV entries. StreamingLLM [12] and LM-infinite [30] identifies the importance of the initial k tokens for generation. H2O [13], ScissorsHand [15] and RoCo [14] recognize crucial KV based on attention scores, while SnapKV [16] utilizes attention within an observation window. Recent works explore improved budget allocation strategies. Pyramid [17, 18] allocates more cache to lower layers, whereas AdaKV [31], HeadKV [19], and DuoAttention [20] focus on inter-head differences. However, eviction causes irreversible information loss, potentially degrading generation quality.
KV Cache Merging. KV cache merging combines less important KV entries instead of permanently discarding them. DMC [22] learns when and how to merge through training, resulting in a lack of generalization and requiring additional processing. In contrast, CaM [21] adaptively merges evicted value states into others but does not merge the corresponding keys. Recently, D2O [23] selects merge candidates and assigns merging weights based on cosine similarity between key states, while KVMerger [24] introduces a variant of the Agglomerative Hierarchical Clustering algorithm to group merge candidates and computes merging weights using Gaussian Kernel Weights. However, these methods fail to maintain attention consistency before and after weighted merging, leading to output perturbation. We propose a novel merging approach designed to eliminate output perturbation, supported by theoretical analysis and extensive comparisons.
3 Methodology
3.1 Preliminary: Inference with KV Cache
We first introduce the attention computation process with KV cache. For simplicity, we consider a single attention head at one layer. Let the attention module’s weight matrices be $W_{q},W_{k},W_{v}\in\mathbb{R}^{d\times d}$, where $d$ denotes the hidden dimension. During the prefill stage, given an input prompt tensor $X_{L}\in\mathbb{R}^{L\times d}=[x_{1},x_{2},\dots,x_{L}]$, where $L$ represents the prompt length, the KV states are computed and stored in the KV cache as follows:
$K_{L}=X_{L}W_{k}=[k_{1},k_{2},\dots,k_{L}],\quad V_{L}=X_{L}W_{v}=[v_{1},v_{2},\dots,v_{L}].$ (1)
In the decoding phase, KV cache are repeatedly utilized, while the newly computed KV pairs are continuously appended to it. Specifically, given the input at the $t$-th generation step, $x_{t}\in\mathbb{R}^{d}$, the KV cache update and attention computation are performed as follows:
$K_{t}=[K_{t-1},k_{t}],V_{t}=[V_{t-1},v_{t}],\quad k_{t}=x_{t}W_{k},v_{t}=x_{t}W_{v},$ (2)
$A^{t}=\text{softmax}\left(\frac{q_{t}K_{t}^{T}}{\sqrt{d}}\right),q_{t}=x_{t}W_{q};\quad o_{t}=\sum_{i=1}^{t}A_{i}^{t}v_{i}=\frac{\sum_{i=1}^{t}s^{t}_{i}v_{i}}{\sum_{i=1}^{t}s^{t}_{i}},s^{t}_{i}=e^{\frac{q_{t}k_{i}}{\sqrt{d}}}$ (3)
KV cache effectively reduces redundant computation, but at the cost of increased memory consumption. Therefore, an important challenge is to compress the KV cache while maintaining model performance.
3.2 Rethinking KV Cache Eviction and Merging
Eviction and merging methods reduce memory usage by decreasing the number of stored KV pairs. The core motivation behind these studies is to minimize the impact of cache compression on the output. A fundamental subtask is to ensure that the output $(o_{t})$ remains as close as possible before and after compression at the current step. However, our analysis shows that existing methods inevitably introduce output perturbation and can not accomplish this task.
Perturbation in KV Cache Eviction. Eviction methods discard KV pairs deemed unimportant. Suppose we discard the pair $(k_{e},v_{e})$, and denote the output as $o^{\prime}_{t}$. Based on Equation 3, we obtain:
$o^{\prime}_{t}=\frac{\sum_{i=1,i\neq e}^{t}s^{t}_{i}v_{i}}{\sum_{i=1,i\neq e}^{t}s^{t}_{i}}=\frac{1}{1-A^{t}_{e}}\left(o_{t}-A^{t}_{e}v_{e}\right).$ (4)
Remark 3.1.
Equation 4 reveals that evicting $(k_{e},v_{e})$ causes $o^{\prime}_{t}$ to deviate from $o_{t}$, with the deviation primarily determined by $A_{e}^{t}$. This formally explains why eviction methods generally prioritize discarding KV pairs with lower attention scores.
Although current methods optimize eviction and cache allocation strategies [18, 19] to minimize output impact, they cannot eliminate the perturbation in Equation 4. Previous studies have indicated that attention is not always sparse, especially in tasks requiring full context, as shown in Figure 3. Moreover, evicted KV may become important later, but irreversible eviction leads to permanent loss.
Figure 3: (a) Cumulative distribution of attention scores. Retaining the top-$k$ tokens does not always preserve the majority of scores. (b) Proportion of to-be-evicted prompt tokens appearing in the top-20% attention scores during generation (compression rate = 20%). (c) Each token’s variance of its attention scores at each generation step (blue dots) is greater than the average variance within a sliding window (orange dots). (d) Relative errors for prediction of KeepKV and existing methods.
Attention Sag in KV Cache Merging. Merging methods integrate less important KV into others rather than discarding them. Existing studies typically use weighted merging [22, 23, 24]; formally, merging $(k_{e},v_{e})$ into $(k_{c},v_{c})$ is expressed as:
$k_{r}=w_{e}k_{e}+w_{c}k_{c},\quad v_{r}=w_{e}v_{e}+w_{c}v_{c}.$ (5)
Here, $(k_{r},v_{r})$ are the merged vectors, with weights $w_{e},w_{c}$ determined by the merging method. In D2O [23], they depend on the cosine similarity between $k_{e}$ and $k_{c}$, while in KVMerger [24], they are computed using Gaussian Kernel values. The weights satisfy the normalization condition $w_{e}+w_{c}=1$. However, this widely used convex combination method also introduces output perturbations:
Theorem 3.2.
Current weighted merging (convex combination) methods reduce the merged KV pair’s attention score compared to the sum of the original scores before merging, i.e., ${A^{\prime}}_{r}^{t}<A_{e}^{t}+A_{c}^{t}$, ultimately leading to $\left\|o^{\prime}_{t}-o_{t}\right\|>0$.
The formal proof is in Appendix A.2. We term this attention inconsistency from merging as Attention Sag and Figure 3 (c) illustrates this phenomenon. We provide an intuitive comprehension: existing methods merge multiple vectors into one, treating it equivalently as any other single vector in subsequent attention computations. This erases merging history, making it impossible to distinguish whether a KV pair is original or has absorbed numerous others.
3.3 Method: KeepKV
3.3.1 Electoral Votes and ZIP-Merging
Electoral Votes. To address Attention Sag, we propose the Electoral Votes mechanism, which records the number of merges $p_{i}$ (initialized to $1$) each KV pair undergoes. A natural analogy is the Electoral College system system, where electors hold votes proportional to their state’s population rather than a uniform share. The attention score of each KV is then scaled by its votes to approximate the original multiple KV’s influence before merging. For example, if a KV pair $(k_{r},v_{r})$ has a vote count of $p_{r}=3$, it is equivalent to three identical and independent instances of $(k_{r},v_{r})$ participating in the attention computation. Formally, the outputs before ($o_{t}$) and after merging ($o^{\prime}_{t}$) are defined as follows:
$o_{t}=\frac{\sum_{i=1}^{t}p_{i}s_{i}^{t}v_{i}}{\sum_{i=1}^{t}p_{i}s_{i}^{t}},\quad o^{\prime}_{t}=\frac{\sum_{i=1,i\neq e,c}^{t}p_{i}s_{i}^{t}v_{i}+p_{r}s_{r}^{t}v_{r}}{\sum_{i=1,i\neq e,c}^{t}p_{i}s_{i}^{t}+p_{r}s_{r}^{t}},p_{r}=p_{e}+p_{c}.$ (6)
Zero Inference-Perturbation Merging (ZIP-Merging). The Electoral Votes mechanism enables the elimination of output perturbations. We define the merging equations and theorem as follows:
$k_{r}=\frac{(w_{e}k_{e}+w_{c}k_{c})\ln{\frac{w_{e}+w_{c}}{p_{e}+p_{c}}}}{w_{e}\ln{s_{e}^{t}}+w_{c}\ln{s_{c}^{t}}},v_{r}=\frac{w_{e}v_{e}+w_{c}v_{c}}{w_{e}+w_{c}},\quad w_{e}=p_{e}s_{e}^{t},w_{c}=p_{c}s_{c}^{t}.$ (7)
Theorem 3.3.
The merging method in Equation 7 is perturbation-free, that is, $\left\|o^{\prime}_{t}-o_{t}\right\|=0$
Remark 3.4.
The proof is in Appendix A.3. Intuitively, our method ensures attention consistency by preserving historical information via Electoral Votes and applying proper scaling (ZIP-Merging) to $(k_{r},v_{r})$ instead of a convex combination.
This theorem confirms that our novel merging approach can eliminate output perturbations and complete the subtask introduced at the beginning of Section 3.2. However, its applicability remains limited to the current iteration, and extending it to multi-step generation requires additional design.
3.3.2 Extending to Multi-Step Generation
EMA Attention Scores. For ZIP-Merging to be effective in real-world multi-step generation, a solid comprehension of attention score dynamics is essential. Fortunately, empirical observations show that attention scores exhibit strong locality (Figure 3 (d)), meaning a token’s attention scores evolve smoothly across adjacent steps, which is also validated by prior studies [18, 21, 25]. From this, we employ the Exponential Moving Average (EMA) [32, 33] with bias correction, a widely used technique in time-series analysis, formulated as follows:
$\hat{s^{t}}=\frac{1}{1-{\alpha}^{t}}S^{t},\quad S^{t}=\left\{\begin{array}[]{lr}\sum_{k=t-w}^{t}(1-\alpha){\alpha}^{t-k}s^{k},\quad t=L&\\ \alpha S^{t-1}+(1-\alpha)s^{t},\quad t>L&\end{array}\right.$ (8)
Note that after the prefill stage, we compute EMA scores using a recent window of length $w$ rather than the entire sequence to obtain a more accurate estimation [16, 18]. We find that this method outperforms mainstream approaches, such as cumulative attention and sliding window averaging, in predicting attention scores. Building on this, replacing all score $s_{i}^{t}$ in Equation 7 with our EMA scores $\hat{s_{i}^{t}}$ from Equation 8 successfully achieves the extension to multi-step generation. Consequently, the future output perturbation becomes estimable and controllable. We present the following theorem and lemma(proof in Appendix A.4):
Theorem 3.5.
For the $t^{\prime}$-th step, let $\left|1-\frac{\hat{s_{i}^{t^{\prime}}}}{s_{i}^{t^{\prime}}}\right|\leq\epsilon,\epsilon<1$, the output perturbation satisfies $\Theta_{t^{\prime}}<\frac{2\epsilon(1+\epsilon)\gamma}{(1-\epsilon)^{2}}$, provided that $\left\|v_{i}-v_{j}\right\|\leq\gamma,\forall i\in[t^{\prime}],j\in\{e,c\}$.
Lemma 3.6.
As the prediction error $\epsilon$ decreases and the merged candidates become increasingly similar, the output perturbation reduces to zero. That is, when either $\epsilon=0$ or $(k_{e},v_{e})=(k_{c},v_{c})$, we have: $\Theta_{t^{\prime}}=0$.
Similarity-driven merging. Lemma 3.6 shows that output perturbation decreases as prediction error $\epsilon$ reduces, and closer merging objects result in lower perturbation. Clearly, if the merged KV pairs are identical, retaining one pair and setting its Electoral Votes to 2 introduces no error in subsequent computations. This provides a theoretical justification for prior merging strategies favoring high-similarity KV pairs [23, 24]. Following this, we merge each evicted KV pair with the retained one having the highest cosine similarity of keys, using a predefined threshold $T$ to determine whether merging should occur, avoiding the overhead of dynamic adjustments like D2O [23].
We present the workflow of KeepKV in Figure 4. Notably, KeepKV imposes no specific constraints on cache allocation or token selection. It can directly integrate with common token selection methods by designating the merging pairs based on their eviction and retention sets, and it is also compatible with various cache allocation strategies. Thus, KeepKV demonstrates strong adaptability and can be combined with a range of mainstream cache compression methods, significantly enhancing both compression capability and generation quality.
Figure 4: Illustrative example of KeepKV. (0) $(k_{e},v_{e})$ is selected for eviction by specific compression method. (1) The retained KV with the highest cosine similarity, $(k_{c},v_{c})$, is selected. (2) EMA attention scores are updated. (3) ZIP-Merging is performed. (4) Consequently, with the Electoral Votes, the compressed KV can preserve the influence of the original KV in attention computations.
4 Experiment
4.1 Experiment Settings
Tasks We evaluate KeepKV on datasets with standard and extended context lengths, covering question-answering, summarization, and synthetic tasks. Specifically, for question-answering, we utilize the COPA [34], MathQA [35], and OpenBookQA [36] tasks from the lm-eval-harness framework [37]. For summarization, we employ the XSUM[38] and CNN/DailyMail[39] tasks provided by the HELM framework. To assess performance on long-context tasks, we adopt LongBench[40], which effectively examines the algorithm’s compression capabilities across diverse subtasks, including single-document QA, multi-document QA, summarization, and synthetic tasks.
Models and baselines. Our evaluation is based on several representative LLMs, including OPT[41], Llama-2 [1], Llama-3 [2], and Mistral [3]. We compare our method against multiple baseline approaches: representative cache eviction methods such as Streaming [12], H2O [13] and PyramidInfer [18], and prominent cache merging methods including CaM [21] and D2O [23]. More detailed comparison results are provided in the Appendix.
Implementation. In our primary experiments, we set the merging threshold $T$ to $0.8$ and the exponential prediction coefficient $\beta=1.2$. For token selection and cache allocation, we follow the strategy recommended by PyramidInfer [18], which allocates fixed cache budgets, making it simple, and efficient. And it is sufficient to demonstrate the advantages of our algorithm. In contrast, D2O [23] applies dynamic allocation based on extra computation after prefill phase for each sequence. We implement KeepKV using the Hugging Face Transformers [42] and conduct all experiments on NVIDIA A100 80GB GPUs.
4.2 Accuracy on KV Cache Compression Ratios
In Figure 5, we benchmark KeepKV on both the lm-eval-harness and HELM frameworks, comparing the fully cached KV version against multiple KV cache compression methods, including our proposed KeepKV. The x-axis represents the compression ratio, defined as the ratio between the compressed KV cache budget and the prompt length $L$. The results demonstrate that KeepKV consistently outperforms all other compression methods across various compression ratios. Particularly at extremely low compression rates, KeepKV achieves significantly better performance, highlighting its superior compression capability to retain maximal information within highly constrained memory budgets while effectively minimizing output perturbations introduced by compression.
Figure 5: Performance of KeepKV and other methods for LLama backbones on summarization datasets including XSUM and CNN/Daily Mail
4.3 Accuracy on Long-context Tasks
We evaluate KeepKV on the LongBench across Llama and Mistral model families . To thoroughly assess compression capabilities on long-context tasks, we set the KV cache budget ratio to $10\%$. The results indicate that KeepKV achieves performance closer to the full-cache baseline on most tasks, maintaining high generation quality despite limited cache availability. Notably, KeepKV significantly outperforms eviction-based methods. Furthermore, KeepKV also surpasses existing KV-cache merging methods, underscoring the effectiveness of our carefully designed merging strategy in enhancing output accuracy.
5 Conclusion
In this paper, we conduct a comprehensive analysis of the impact of KV cache compression on attention computation and propose KeepKV, which introduces the Electoral Votes mechanism and Zero Inference-Perturbation Merging to adaptively and dynamically merge the KV cache while minimizing output disturbance. KeepKV effectively preserves more information within limited memory, significantly mitigating the adverse effects of KV cache compression on generation quality. Our experiments demonstrate that KeepKV achieves performance closest to that of the full cache across various compression ratios. It also excels in both standard and long-context tasks. We believe KeepKV provides a novel perspective and a powerful tool for advancing KV cache compression methods, laying the foundation for efficient LLM inference.
6 Acknowledgement
We are grateful to Chenhong He, Ruijie Miao, Yuhan Wu and Yanshu Wang from Peking University for their insightful discussions and helpful suggestions throughout the development of this research. We thank ByteDance Ltd. for providing technical support during the internship period.
Appendix A Theoretical Analysis
Recently, many studies have analyzed KV cache compression strategies in LLM inference from a theoretical perspective [13, 15, 16, 18, 21, 23, 24, 43]. Overall, the primary objective of most existing works can be summarized as minimizing the impact of compression on the output. For instance, existing eviction-based methods and cache allocation strategies [13, 15, 14, 16, 17, 18, 19, 20, 31] all aim to maximize the retention of essential information within limited memory by evicting less important tokens or reducing cache allocation in non-critical heads and layers based on empirical observations of attention distributions. However, eviction inevitably leads to irreversible information loss, which has motivated the development of KV cache merging methods [21, 22, 23, 24]. Despite this, key challenges such as the selection of merging candidates and the assignment of merging weights remain largely unexplored, with a lack of systematic theoretical foundations. In this work, we introduce a novel perspective distinct from prior approaches. We formulate the problem as eliminating output perturbation and derive a novel merging method by analyzing the attention computation process. First, we introduce Electoral Votes mechanism, making the elimination of output perturbation feasible. Then, we derive a merging computation formula to eliminate perturbation at the current step. Finally, we extend this framework to multi-step generation, providing a theoretical guarantee for output perturbation and offering a reasonable explanation for mainstream similarity-based merging candidates selection methods.
Specifically, in Section A.1, we demonstrate the unavoidable output perturbation caused by KV cache eviction. In Section A.2, we discuss the attention sag issue in existing KV cache merging methods and provide a formal proof. In Section A.3, we present the derivation process of our KeepKV merging method. Finally, in Section A.4, we provide a theoretical guarantee for the output perturbation of KeepKV, including proofs for Theorem 3.5 and its associated lemma. The symbolic representation of attention computation process remains consistent with Section 3.
A.1 Perturbation in KV Cache Eviction
Eviction methods discard KV pairs deemed unimportant. We denote the first generation step in the decoding phase as the $(L+1)$-th generation step, where $L$ represents the prompt length. And for a positive integer n, let $[n]:=\{1,2,...,n\}$. At $t$-th generation step, let $K_{e}=\{e_{1}e_{2},...,e_{m}\},m\in[t]$ denoted the index of to-be-evicted cache. Based on Equation 3, the output after eviction $(o^{\prime}_{t})$ is:
$o^{\prime}_{t}=\sum_{i=1,i\notin K_{e}}^{t}A_{i}^{{}^{\prime}t}v_{i},\quad A_{i}^{{}^{\prime}t}=\frac{s^{t}_{i}}{\sum_{i=1,i\notin K_{e}}^{t}s^{t}_{i}}.$ (9)
By transforming $o^{\prime}_{t}$ towards $o_{t}$, we obtain:
$o^{\prime}_{t}$ $=\frac{\sum_{i=1}^{t}s^{t}_{i}}{\sum_{i=1}^{t}s^{t}_{i}-\sum_{j\in K_{e}}^{t}s^{t}_{j}}*\frac{\sum_{i=1}^{t}s^{t}_{i}v_{i}-\sum_{j\in K_{e}}^{t}s^{t}_{j}v_{j}}{\sum_{i=1}^{t}s^{t}_{i}}$ (10)
$=\frac{\sum_{i=1}^{t}s^{t}_{i}}{\sum_{i=1}^{t}s^{t}_{i}-\sum_{j\in K_{e}}^{t}s^{t}_{j}}\left(o_{t}-\frac{\sum_{j\in K_{e}}^{t}s^{t}_{j}v_{j}}{\sum_{i=1}^{t}s^{t}_{i}}\right)$
$=\frac{\sum_{i=1}^{t}A^{t}_{i}}{\sum_{i=1}^{t}A^{t}_{i}-\sum_{j\in K_{e}}^{t}A^{t}_{j}}\left(o_{t}-\frac{\sum_{j\in K_{e}}^{t}A^{t}_{j}v_{j}}{\sum_{i=1}^{t}A^{t}_{i}}\right)$
$=\frac{1}{1-\sum_{j\in K_{e}}^{t}A^{t}_{j}}\left(o_{t}-\sum_{j\in K_{e}}^{t}A^{t}_{j}v_{j}\right).$
Equation 10 indicates that the difference between $o^{\prime}_{t}$ and $o_{t}$ decreases as the attention score of the evicted KV $(A^{t}_{j},j\in K_{e})$ diminishes. When the total score of $K_{e}$ becomes negligible, the output perturbation at the current step approaches zero. This formally explains why eviction methods generally prioritize discarding KV pairs with lower attention scores.
However, existing studies [44] have shown that attention can be relatively dispersed in certain tasks, meaning that evicting even a small number of tokens can have a non-negligible impact. Furthermore, as the compression ratio increases, evicted tokens will account for a significant portion of the attention scores, exacerbating the degradation of generation quality.
A.2 Attention Sag in KV Cache Merging
Merging methods integrate less important KV into others rather than discarding them directly. Specifically, mainstream studies select, for each KV pair to be evicted, a merging target among the preserved KVs, allowing many-to-one merges. Typically, weighted merging rather than direct averaging is used, with weights satisfying a normalization constraint, i.e., the merged vectors are obtained via convex combinations. Formally, merging the evicted pairs $(k_{j},v_{j}),j\in K_{e}$ into a preserved pair $(k_{c},v_{c})$ yields a new KV pair $(k_{r},v_{r})$, defined as follows:
$k_{r}=w_{c}k_{c}+\sum_{j\in K_{e}}w_{j}k_{j},v_{r}=w_{c}v_{c}+\sum_{j\in K_{e}}w_{j}v_{j},\quad w_{c}+\sum_{j\in K_{e}}w_{j}=1$ (11)
Let $K^{\prime}_{e}=K_{e}\cup\{c\}$, representing the index of the original KVs before merging. For instance, the weight $w_{j}$ in D2O [23] is computed based on the cosine similarity between key vectors, whereas for KVmerger [24], it is calculated based on the Gaussian Kernel value. Formally, these are represented as follows:
$\ w_{j_{D2O}}=\frac{\exp(\cos\theta_{\mathbf{k_{j}},\mathbf{k_{c}}})}{\sum_{j\in K^{\prime}_{e}}\exp(\cos\theta_{\mathbf{k_{j}},\mathbf{k_{c}}})};\quad w_{j_{KVMerger}}=\frac{\exp(-\frac{||k_{j}-k_{c}||^{2}}{2\sigma^{2}})}{\sum_{j\in K^{\prime}_{e}}\exp(-\frac{||k_{j}-k_{c}||^{2}}{2\sigma^{2}})}.$ (12)
However, the widely adopted convex combination approach also introduces output disturbances, as stated in the following theorem:
Theorem A.1 (Formal version of Theorem 3.2).
The merging method indicated by Equation 11 causes the attention score of the merged KV to become less than the sum of attention scores from the original multiple KVs merged into it, independently of the specific weighting scheme. Formally, this implies: ${A^{\prime}}_{r}^{t}<\sum_{j\in K^{\prime}_{e}}A^{t}_{j}$, ultimately leading to: $\left\|o^{\prime}_{t}-o_{t}\right\|>0$.
Proof.
The attention score and output after merging can be expressed as:
$o^{\prime}_{t}=\sum_{i=1,i\notin K^{\prime}_{e}}^{t}{A^{\prime}}_{i}^{t}v_{i}+{A^{\prime}}_{r}^{t}v_{r}=\frac{\sum_{i=1,i\notin K^{\prime}_{e}}^{t}s^{t}_{i}v_{i}+s^{t}_{r}v_{r}}{\sum_{i=1,i\notin K^{\prime}_{e}}^{t}s^{t}_{i}+s^{t}_{r}}.$ (13)
First, we compare the denominators of $A^{t}$ and $A^{\prime t}$, formally proving $\sum_{i=1,i\notin K^{\prime}_{e}}^{t}s^{t}_{i}+s^{t}_{r}<\sum_{i=1}^{t}s^{t}_{i}$:
$\sum_{i=1,i\notin K^{\prime}_{e}}^{t}s^{t}_{i}+s^{t}_{r}=\sum_{i=1}^{t}s^{t}_{i}+s^{t}_{r}-\sum_{i\in K^{\prime}_{e}}s^{t}_{i}=e^{\frac{q^{t}k_{r}}{\sqrt{d}}}-\sum_{i\in K^{\prime}_{e}}e^{\frac{q^{t}k_{i}}{\sqrt{d}}}$ (14)
Substituting Equation 11, and applying the Weighted AM–GM Inequality, we have:
$e^{\frac{q^{t}k_{r}}{\sqrt{d}}}-\sum_{i\in K^{\prime}_{e}}e^{\frac{q^{t}k_{i}}{\sqrt{d}}}$ $=e^{\frac{q^{t}(\sum_{i\in K^{\prime}_{e}}w_{i}k_{i})}{\sqrt{d}}}-\sum_{i\in K^{\prime}_{e}}e^{\frac{q^{t}k_{i}}{\sqrt{d}}}$ (15)
$=\prod_{i\in K^{\prime}_{e}}{(e^{\frac{q^{t}k_{i}}{\sqrt{d}}})}^{w_{i}}-\sum_{i\in K^{\prime}_{e}}e^{\frac{q^{t}k_{i}}{\sqrt{d}}}$
$\leq\sum_{i\in K^{\prime}_{e}}w_{i}e^{\frac{q^{t}k_{i}}{\sqrt{d}}}-\sum_{i\in K^{\prime}_{e}}e^{\frac{q^{t}k_{i}}{\sqrt{d}}}<0$
Thus,
$\sum_{i=1,i\notin K^{\prime}_{e}}^{t}s^{t}_{i}+s^{t}_{r}=\sum_{i=1}^{t}s^{t}_{i}+(s^{t}_{r}-\sum_{i\in K^{\prime}_{e}}s^{t}_{i})<\sum_{i=1}^{t}s^{t}_{i}$ (16)
Since the sum of the normalized attention scores equals one, and given that $\sum_{i=1,i\notin K^{\prime}_{e}}^{t}s^{t}_{i}+s^{t}_{r}<\sum_{i=1}^{t}s^{t}_{i}$, we obtain:
${A^{\prime}}_{r}^{t}=1-\frac{\sum_{i=1,i\notin K^{\prime}_{e}}^{t}s^{t}_{i}}{\sum_{i=1,i\notin K^{\prime}_{e}}^{t}s^{t}_{i}+s^{t}_{r}}<1-\frac{\sum_{i=1,i\notin K^{\prime}_{e}}^{t}s^{t}_{i}}{\sum_{i=1}^{t}s^{t}_{i}}=\sum_{i\in K^{\prime}_{e}}{A}_{i}^{t}$ (17)
Similarly, we can derive:
${A^{\prime}}_{j}^{t}=\frac{s^{t}_{j}}{\sum_{i=1,i\notin K^{\prime}_{e}}^{t}s^{t}_{i}+s^{t}_{r}}>\frac{s^{t}_{j}}{\sum_{i=1}^{t}s^{t}_{i}}={A}_{j}^{t},\quad j\neq r$ (18)
Finally, the output perturbation can be represented as:
$\left\|o^{\prime}_{t}-o_{t}\right\|$ $=\left\|(\sum_{i=1,i\notin K^{\prime}_{e}}^{t}A_{i}^{{}^{\prime}t}v_{i}+{A^{\prime}}_{r}^{t}v_{r})-\sum_{i=1}^{t}A_{i}^{t}v_{i}\right\|$ (19)
$=\left\|\sum_{i=1,i\notin K^{\prime}_{e}}^{t}(A_{i}^{{}^{\prime}t}-A_{i}^{t})v_{i}+({A^{\prime}}_{r}^{t}v_{r}-\sum_{i\in K^{\prime}_{e}}{A}_{i}^{t}v_{i})\right\|$
$=\left\|\sum_{i=1,i\notin K^{\prime}_{e}}^{t}(A_{i}^{{}^{\prime}t}-A_{i}^{t})v_{i}+\sum_{i\in K^{\prime}_{e}}(w_{i}{A^{\prime}}_{r}^{t}-A_{i}^{t})v_{i}\right\|$
In the above expression, all vector coefficients are nonzero. Moreover, due to the high dimensionality and sparsity of the KV cache [24, 43], the vectors are almost linearly independent. In practical inference scenarios, it is nearly impossible for them to form a zero vector through linear combination. Consequently, we have: $\left\|o^{\prime}_{t}-o_{t}\right\|>0$. ∎
We term this phenomenon as Attention Sag, indicating that improper merging methods result in a reduced attention score for the newly merged vector, while attention scores of unmerged KVs relatively increase. This leads to output disturbances and ultimately degrades generation quality.
A.3 KeepKV Merging Method
In Section 3.3.1, we introduced the concept of merging count via the Electoral Votes mechanism, aiming for a KV pair with vote count $p_{i}$ to be equivalent, in attention computation, to $p_{i}$ independent occurrences of this KV. Moreover, the vote count of the merged KV equals the sum of vote counts before merging. Formally, the outputs before $(o_{t})$ and after merging $(o^{\prime}_{t})$ can be expressed as follows:
$o_{t}=\frac{\sum_{i=1}^{t}p_{i}s_{i}^{t}v_{i}}{\sum_{i=1}^{t}p_{i}s_{i}^{t}},\quad o^{\prime}_{t}=\frac{\sum_{i=1,i\notin K^{\prime}_{e}}^{t}p_{i}s_{i}^{t}v_{i}+p_{r}s_{r}^{t}v_{r}}{\sum_{i=1,i\notin K^{\prime}_{e}}^{t}p_{i}s_{i}^{t}+p_{r}s_{r}^{t}},p_{r}=\sum_{i\in K^{\prime}_{e}}p_{i}.$ (20)
Next, we demonstrate how our new merging approach in Section 3.3.1 can be derived naturally from the objective of eliminating output disturbances, which consequently serves as a direct proof for Theorem 3.3.
Based on Equation 20, setting $\left\|o^{\prime}_{t}-o_{t}\right\|=0$, we obtain:
$\sum_{i=1}^{t}p_{i}s_{i}^{t}v_{i}=\sum_{i=1,i\notin K^{\prime}_{e}}^{t}p_{i}s_{i}^{t}v_{i}+p_{r}s_{r}^{t}v_{r},\quad\sum_{i=1}^{t}p_{i}s_{i}^{t}=\sum_{i=1,i\notin K^{\prime}_{e}}^{t}p_{i}s_{i}^{t}+p_{r}s_{r}^{t}$ (21)
which implies:
$\sum_{i\in K^{\prime}_{e}}p_{i}s_{i}^{t}v_{i}=p_{r}s_{r}^{t}v_{r},\quad\sum_{i\in K^{\prime}_{e}}p_{i}s_{i}^{t}=p_{r}s_{r}^{t}$ (22)
Dividing the two expressions above, we obtain the expression of $v_{r}$:
$v_{r}=\frac{\sum_{i\in K^{\prime}_{e}}p_{i}s_{i}^{t}v_{i}}{\sum_{i\in K^{\prime}_{e}}p_{i}s_{i}^{t}}$ (23)
Similarly, let $k_{r}=C(\sum_{i\in K^{\prime}_{e}}p_{i}s_{i}^{t}k_{i})$, substituting this into $\sum_{i\in K^{\prime}_{e}}^{t}p_{i}s_{i}^{t}=p_{r}s_{r}^{t}$ from Equation 22 and solving, we obtain:
$C=\frac{\ln{\frac{\sum_{i\in K^{\prime}_{e}}p_{i}s_{i}^{t}}{\sum_{i\in K^{\prime}_{e}}p_{i}}}}{\sum_{i\in K^{\prime}_{e}}p_{i}s_{i}^{t}\ln{s_{i}^{t}}}$ (24)
Finally, we derive the merging expression:
$k_{r}=\frac{(\sum_{i\in K^{\prime}_{e}}p_{i}s_{i}^{t}k_{i})\ln{\frac{\sum_{i\in K^{\prime}_{e}}p_{i}s_{i}^{t}}{\sum_{i\in K^{\prime}_{e}}p_{i}}}}{\sum_{i\in K^{\prime}_{e}}p_{i}s_{i}^{t}\ln{s_{i}^{t}}},\quad v_{r}=\frac{\sum_{i\in K^{\prime}_{e}}p_{i}s_{i}^{t}v_{i}}{\sum_{i\in K^{\prime}_{e}}p_{i}s_{i}^{t}},\quad p_{r}=\sum_{i\in K^{\prime}_{e}}p_{i}$ (25)
Consequently, merging in this manner eliminates the output disturbance in the $t$-step, satisfying: $\left\|o^{\prime}_{t}-o_{t}\right\|=0$. By setting the merging candidates $K^{\prime}_{e}=\{e\}\cup\{c\}$, we obtain Theorem 3.3.
A.4 Error Bound Analysis
After extending KeepKV to multi-step generation in Section 3.3.2, for $t^{\prime}$-step, all $s_{i}^{t}$ terms in Equation 25 are replaced with $\hat{s_{i}^{t^{\prime}}}$ , which represents our estimation of future attention score trends obtained through a certain method. In this case, the merging expressions become:
$k_{r}=\frac{(\sum_{i\in K^{\prime}_{e}}p_{i}\hat{s_{i}^{t^{\prime}}}k_{i})\ln{\frac{\sum_{i\in K^{\prime}_{e}}p_{i}\hat{s_{i}^{t^{\prime}}}}{\sum_{i\in K^{\prime}_{e}}p_{i}}}}{\sum_{i\in K^{\prime}_{e}}p_{i}\hat{s_{i}^{t^{\prime}}}\ln{\hat{s_{i}^{t^{\prime}}}}},\quad v_{r}=\frac{\sum_{i\in K^{\prime}_{e}}p_{i}\hat{s_{i}^{t^{\prime}}}v_{i}}{\sum_{i\in K^{\prime}_{e}}p_{i}\hat{s_{i}^{t^{\prime}}}},\quad p_{r}=\sum_{i\in K^{\prime}_{e}}p_{i}$ (26)
and they satisfy:
$\sum_{i\in K^{\prime}_{e}}p_{i}\hat{s_{i}^{t^{\prime}}}v_{i}=p_{r}\hat{s_{r}^{t}}v_{r},\quad\sum_{i\in K^{\prime}_{e}}p_{i}\hat{s_{i}^{t}}=p_{r}\hat{s_{r}^{t}}$ (27)
Then the perturbation at step $s^{\prime}$, can be expressed as:
$\Theta_{t^{\prime}}$ $=\left\|o_{t^{\prime}}-o^{\prime}_{t^{\prime}}\right\|$ (28)
$=\left\|\frac{\sum_{i=1}^{t^{\prime}}p_{i}s_{i}^{t^{\prime}}v_{i}}{\sum_{i=1}^{t^{\prime}}p_{i}s_{i}^{t^{\prime}}}-\frac{\sum_{i=1,i\notin K^{\prime}_{e}}^{t^{\prime}}p_{i}s_{i}^{t^{\prime}}v_{i}+p_{r}s_{r}^{t^{\prime}}v_{r}}{\sum_{i=1,i\notin K^{\prime}_{e}}^{t^{\prime}}p_{i}s_{i}^{t^{\prime}}+p_{r}s_{r}^{t^{\prime}}}\right\|$
$=\frac{\left\|\sum_{i=1}^{t^{\prime}}p_{i}s_{i}^{t^{\prime}}[\sum_{j\in K^{\prime}_{e}}p_{j}s_{j}^{t^{\prime}}(v_{j}-v_{i})-p_{r}s_{r}^{t^{\prime}}(v_{r}-v_{i})]\right\|}{(\sum_{i=1}^{t^{\prime}}p_{i}s_{i}^{t^{\prime}})(\sum_{i=1,i\notin K^{\prime}_{e}}^{t^{\prime}}p_{i}s_{i}^{t^{\prime}}+p_{r}s_{r}^{t^{\prime}})}$
Substituting the expression for $v_{r}$ from Equation 26 and $\sum_{j\in K^{\prime}_{e}}p_{j}\hat{s_{j}^{t^{\prime}}}=p_{r}\hat{s_{r}^{t^{\prime}}}$ from Equation 27 into the above: s
$\sum_{i=1}^{t^{\prime}}p_{i}s_{i}^{t^{\prime}}[\sum_{j\in K^{\prime}_{e}}p_{j}s_{j}^{t^{\prime}}(v_{j}-v_{i})-p_{r}s_{r}^{t^{\prime}}(v_{r}-v_{i})]$ (29)
$=\sum_{i=1}^{t^{\prime}}p_{i}s_{i}^{t^{\prime}}[\sum_{j\in K^{\prime}_{e}}p_{j}s_{j}^{t^{\prime}}(v_{j}-v_{i})-p_{r}s_{r}^{t^{\prime}}(\frac{\sum_{k\in K^{\prime}_{e}}p_{k}\hat{s_{k}^{t^{\prime}}}v_{k}}{\sum_{k\in K^{\prime}_{e}}p_{k}\hat{s_{k}^{t^{\prime}}}}-v_{i})]$
$=\sum_{i=1}^{t^{\prime}}p_{i}s_{i}^{t^{\prime}}[\sum_{j\in K^{\prime}_{e}}p_{j}s_{j}^{t^{\prime}}(1-\frac{s_{r}^{t^{\prime}}}{\hat{s_{r}^{t^{\prime}}}}\frac{\hat{s_{j}^{t^{\prime}}}}{s_{j}^{t^{\prime}}})(v_{j}-v_{i})]$
Let $\left|1-\frac{\hat{s_{i}^{t^{\prime}}}}{s_{i}^{t^{\prime}}}\right|\leq\epsilon,\epsilon<1$, then $1-\epsilon\leq\frac{\hat{s_{i}^{t^{\prime}}}}{s_{i}^{t^{\prime}}}\leq 1+\epsilon$, thus:
$\left|1-\frac{s_{r}^{t^{\prime}}}{\hat{s_{r}^{t^{\prime}}}}\frac{\hat{s_{j}^{t^{\prime}}}}{s_{j}^{t^{\prime}}}\right|=\left|\frac{\frac{\hat{s_{r}^{t^{\prime}}}}{s_{r}^{t^{\prime}}}-\frac{\hat{s_{j}^{t^{\prime}}}}{s_{j}^{t^{\prime}}}}{\frac{\hat{s_{r}^{t^{\prime}}}}{s_{r}^{t^{\prime}}}}\right|\leq\frac{2\epsilon}{1-\epsilon},\quad j\in K^{\prime}_{e}$ (30)
Let $\left\|v_{j}-v_{i}\right\|\leq\gamma,\forall i\in[t^{\prime}],j\in K^{\prime}_{e}$, where $\gamma$ represents the inherent variation in the input, which cannot be eliminated through algorithmic design. Then, applying the triangle inequality, we obtain:
$\left\|\sum_{i=1}^{t^{\prime}}p_{i}s_{i}^{t^{\prime}}[\sum_{j\in K^{\prime}_{e}}p_{j}s_{j}^{t^{\prime}}(1-\frac{s_{r}^{t^{\prime}}}{\hat{s_{r}^{t^{\prime}}}}\frac{\hat{s_{j}^{t^{\prime}}}}{s_{j}^{t^{\prime}}})(v_{j}-v_{i})]\right\|$ $\leq\sum_{i=1}^{t^{\prime}}p_{i}s_{i}^{t^{\prime}}(\sum_{j\in K^{\prime}_{e}}p_{j}s_{j}^{t^{\prime}}\left|1-\frac{s_{r}^{t^{\prime}}}{\hat{s_{r}^{t^{\prime}}}}\frac{\hat{s_{j}^{t^{\prime}}}}{s_{j}^{t^{\prime}}}\right|\left\|v_{j}-v_{i}\right\|)$ (31)
$\leq\frac{2\epsilon\gamma}{1-\epsilon}(\sum_{i=1}^{t^{\prime}}p_{i}s_{i}^{t^{\prime}})(\sum_{j\in K^{\prime}_{e}}p_{j}s_{j}^{t^{\prime}})$
Substituting this inequality into Equation 28, we obtain:
$\Theta_{t^{\prime}}$ $\leq\frac{2\epsilon\gamma}{1-\epsilon}\frac{(\sum_{i=1}^{t^{\prime}}p_{i}s_{i}^{t^{\prime}})(\sum_{j\in K^{\prime}_{e}}p_{j}s_{j}^{t^{\prime}})}{(\sum_{i=1}^{t^{\prime}}p_{i}s_{i}^{t^{\prime}})(\sum_{i=1,i\notin K^{\prime}_{e}}^{t^{\prime}}p_{i}s_{i}^{t^{\prime}}+p_{r}s_{r}^{t^{\prime}})}$ (32)
$=\frac{2\epsilon\gamma}{1-\epsilon}\frac{\sum_{j\in K^{\prime}_{e}}p_{j}s_{j}^{t^{\prime}}}{\sum_{i=1,i\notin K^{\prime}_{e}}^{t^{\prime}}p_{i}s_{i}^{t^{\prime}}+p_{r}s_{r}^{t^{\prime}}}$
$<\frac{2\epsilon\gamma}{1-\epsilon}\frac{\sum_{j\in K^{\prime}_{e}}p_{j}s_{j}^{t^{\prime}}}{p_{r}s_{r}^{t^{\prime}}}$
Due to Equation 27, we have $\frac{\sum_{j\in K^{\prime}_{e}}p_{j}\hat{s_{j}^{t^{\prime}}}}{p_{r}\hat{s_{r}^{t^{\prime}}}}=1$, then:
$\frac{\sum_{j\in K^{\prime}_{e}}p_{j}s_{j}^{t^{\prime}}}{p_{r}s_{r}^{t^{\prime}}}\leq\frac{\frac{1}{1-\epsilon}\sum_{j\in K^{\prime}_{e}}p_{j}\hat{s_{j}^{t^{\prime}}}}{\frac{1}{1+\epsilon}p_{r}\hat{s_{r}^{t^{\prime}}}}=\frac{1+\epsilon}{1-\epsilon}$ (33)
Thus,
$\Theta_{t^{\prime}}$ $<\frac{2\epsilon\gamma}{1-\epsilon}\frac{\sum_{j\in K^{\prime}_{e}}p_{j}s_{j}^{t^{\prime}}}{p_{r}s_{r}^{t^{\prime}}}<\frac{2\epsilon(1+\epsilon)\gamma}{(1-\epsilon)^{2}}$ (34)
Finally, we obtain the following theorem:
Theorem A.2.
For the $t^{\prime}$-th step, let $\left|1-\frac{\hat{s_{i}^{t^{\prime}}}}{s_{i}^{t^{\prime}}}\right|\leq\epsilon,\epsilon<1$, the output perturbation satisfies $\Theta_{t^{\prime}}<\frac{2\epsilon(1+\epsilon)\gamma}{(1-\epsilon)^{2}}$, provided that $\left\|v_{j}-v_{i}\right\|\leq\gamma,\forall i\in[t^{\prime}],j\in K^{\prime}_{e}$.
Next, we prove the following lemma:
Lemma A.3.
As the prediction error $\epsilon$ decreases and the merged candidates become increasingly similar, the output disturbance reduces to zero. That is, when either $\epsilon=0$ or $(k_{i},v_{i})=(k_{j},v_{j}),\forall i,j\in K^{\prime}_{e}$, we have: $\Theta_{t^{\prime}}=0$.
Proof.
By Theorem A.2, it is easy to obtain that when $\epsilon=0$, $\Theta_{t^{\prime}}<\frac{2\epsilon(1+\epsilon)\gamma}{(1-\epsilon)^{2}}=0$. Next, we prove that when $(k_{i},v_{i})=(k_{j},v_{j}),\forall i,j\in K^{\prime}_{e}$, it also holds that $\Theta_{t^{\prime}}=0$. First, we further expand $(1-\frac{s_{r}^{t^{\prime}}}{\hat{s_{r}^{t^{\prime}}}}\frac{\hat{s_{j}^{t^{\prime}}}}{s_{j}^{t^{\prime}}}),j\in K^{\prime}_{e}$ in Equation 29 by applying Equation 27:
$1-\frac{s_{r}^{t^{\prime}}}{\hat{s_{r}^{t^{\prime}}}}\frac{\hat{s_{j}^{t^{\prime}}}}{s_{j}^{t^{\prime}}}$ $=1-\frac{e^{\frac{q_{t^{\prime}}k_{r}}{\sqrt{d}}}}{\frac{\sum_{i\in K^{\prime}_{e}}p_{i}\hat{s_{i}^{t^{\prime}}}}{\sum_{i\in K^{\prime}_{e}}p_{i}}}\frac{\hat{s_{j}^{t^{\prime}}}}{s_{j}^{t^{\prime}}},\quad j\in K^{\prime}_{e}$ (35)
Substituting the expression for $k_{r}$ from Equation 26 into the above:
$1-\frac{e^{\frac{q_{t^{\prime}}k_{r}}{\sqrt{d}}}}{\frac{\sum_{i\in K^{\prime}_{e}}p_{i}\hat{s_{i}^{t}}}{\sum_{i\in K^{\prime}_{e}}p_{i}}}\frac{\hat{s_{j}^{t^{\prime}}}}{s_{j}^{t^{\prime}}}$ $=1-\frac{{(\prod_{i\in K^{\prime}_{e}}{s_{i}^{t^{\prime}}}^{p_{i}\hat{s_{i}^{t^{\prime}}}})^{\frac{\ln{\frac{\sum_{i\in K^{\prime}_{e}}p_{i}\hat{s_{i}^{t^{\prime}}}}{\sum_{i\in K^{\prime}_{e}}p_{i}}}}{\sum_{i\in K^{\prime}_{e}}p_{i}\hat{s_{i}^{t^{\prime}}}\ln{\hat{s_{i}^{t^{\prime}}}}}}}}{\frac{\sum_{i\in K^{\prime}_{e}}p_{i}\hat{s_{i}^{t^{\prime}}}}{\sum_{i\in K^{\prime}_{e}}p_{i}}}\frac{\hat{s_{j}^{t^{\prime}}}}{s_{j}^{t^{\prime}}}$ (36)
When $(k_{i},v_{i})=(k_{j},v_{j}),\forall i,j\in K^{\prime}_{e}$, it follows that $\forall i\in K^{\prime}_{e},s_{i}^{t^{\prime}}=s^{t^{\prime}},\hat{s_{i}^{t^{\prime}}}=\hat{s^{t^{\prime}}}$, thereby:
$1-\frac{{(\prod_{i\in K^{\prime}_{e}}{s_{i}^{t^{\prime}}}^{p_{i}\hat{s_{i}^{t^{\prime}}}})^{\frac{\ln{\frac{\sum_{i\in K^{\prime}_{e}}p_{i}\hat{s_{i}^{t^{\prime}}}}{\sum_{i\in K^{\prime}_{e}}p_{i}}}}{\sum_{i\in K^{\prime}_{e}}p_{i}\hat{s_{i}^{t^{\prime}}}\ln{\hat{s_{i}^{t^{\prime}}}}}}}}{\frac{\sum_{i\in K^{\prime}_{e}}p_{i}\hat{s_{i}^{t^{\prime}}}}{\sum_{i\in K^{\prime}_{e}}p_{i}}}\frac{\hat{s_{j}^{t^{\prime}}}}{s_{j}^{t^{\prime}}}$ $=1-\frac{{(\prod_{i\in K^{\prime}_{e}}{s^{t^{\prime}}}^{p_{i}\hat{s^{t^{\prime}}}})^{\frac{\ln{\frac{\sum_{i\in K^{\prime}_{e}}p_{i}\hat{s^{t^{\prime}}}}{\sum_{i\in K^{\prime}_{e}}p_{i}}}}{\sum_{i\in K^{\prime}_{e}}p_{i}\hat{s^{t^{\prime}}}\ln{\hat{s^{t^{\prime}}}}}}}}{\frac{\sum_{i\in K^{\prime}_{e}}p_{i}\hat{s^{t^{\prime}}}}{\sum_{i\in K^{\prime}_{e}}p_{i}}}\frac{\hat{s^{t^{\prime}}}}{s^{t^{\prime}}}$ (37)
$=1-\frac{s^{t^{\prime}}}{\hat{s^{t^{\prime}}}}\frac{\hat{s^{t^{\prime}}}}{s^{t^{\prime}}}=0$
Under this condition, it follows that Equation 29 equals 0, i.e.,
$\sum_{i=1}^{t^{\prime}}p_{i}s_{i}^{t^{\prime}}[\sum_{j\in K^{\prime}_{e}}p_{j}s_{j}^{t^{\prime}}(v_{j}-v_{i})-p_{r}s_{r}^{t^{\prime}}(v_{r}-v_{i})]=\sum_{i=1}^{t^{\prime}}p_{i}s_{i}^{t^{\prime}}[\sum_{j\in K^{\prime}_{e}}p_{j}s_{j}^{t^{\prime}}(1-\frac{s_{r}^{t^{\prime}}}{\hat{s_{r}^{t^{\prime}}}}\frac{\hat{s_{j}^{t^{\prime}}}}{s_{j}^{t^{\prime}}})(v_{j}-v_{i})]=0$ (38)
Finally, Substituting it into Equation 28, we obtain $\Theta_{t^{\prime}}=0$. ∎
Remark A.4.
This lemma provides a theoretical justification for prior merging strategies favoring high-similarity KV pairs [23, 24]. Meanwhile, we offer an intuitive interpretation: if the merged two KV pairs are identical, i.e., $(k_{e},v_{e})=(k_{c},v_{c})$, retaining one pair and setting its Electoral Votes to 2 introduces no error in subsequent computations.
In this section, we have proven Theorem A.2 and Lemma A.3, which provide guarantees on the output perturbation in multi-step generation—an aspect that existing methods struggle to achieve. Moreover, our method demonstrates superior performance across various experimental evaluations. However, it should be acknowledged that predicting attention distributions further into the future is inherently challenging, leading to a significant increase in the estimated perturbation upper bound. Furthermore, the inherent input differences $\gamma$ cannot be ignored, representing a fundamental problem in KV cache compression—namely, the inability to perfectly compress the KV cache into a smaller memory without any loss of information. Nevertheless, our work introduces a new perspective and analytical approach to studying KV cache eviction and merging algorithms, which we hope will inspire future research.